# Geometric Calibration and Neutral Zones for Uncertainty-Aware Multi-Class Classification

### v1

@@ Abstract
Modern artificial intelligence systems make critical decisions yet often fail silently when uncertain. We develop a geometric framework for post-hoc calibration of neural network probability outputs, treating probability vectors as points on the (c1)(c-1)-dimensional probability simplex equipped with the Fisher–Rao metric. Our approach yields Additive Log-Ratio (ALR) calibration maps that reduce exactly to Platt scaling for binary problems (Proposition 1) while extending naturally to multi-class settings—providing a principled generalization that existing methods lack. Complementing calibration, we define geometric reliability scores based on Fisher–Rao distance and construct neutral zones for principled deferral of uncertain predictions.
Theoretical contributions include: (i) consistency of the calibration estimator at rate $O_{p}(n^{-1/2})$ via M-estimation theory (Theorem 1), and (ii) tight concentration bounds for reliability scores with explicit sub-Gaussian parameters enabling sample size calculations for validation set design (Theorem 2). We conjecture Neyman–Pearson optimality of our neutral zone construction based on connections to Bhattacharyya coefficients. Empirical validation on Adeno-Associated Virus classification demonstrates that the two-stage framework (calibration followed by reliability-based deferral) captures 72.5% of errors while deferring 34.5% of samples. Notably, this operational gain is achievable with any well-calibrated probability output; the contribution of geometric calibration lies in its theoretical foundations rather than empirical superiority over simpler alternatives. This work bridges information geometry and statistical learning, offering formal guarantees relevant to applications requiring rigorous validation.
@@
Preprint Note: This manuscript is under preparation for journal submission. Section 5 presents open problems where we invite community collaboration. Comments welcome at soumojit.das@wsu.edu.
@@ 1 Introduction
The fundamental issue is miscalibration: neural network probability outputs do not reflect true likelihood of correctness. Guo et al. [1] demonstrated that modern deep neural networks exhibit Expected Calibration Error (ECE) often exceeding 10% on standard benchmarks. This manifests as systematic overconfidence, where networks assign high probabilities to misclassified samples [2], and complete absence of uncertainty awareness [3].
@@ 1.3 Our Approach and Contributions
a Riemannian manifold equipped with the Fisher-Rao metric (Definition 1). This geometric perspective reveals that existing calibration methods—which apply transformations treating probability vectors as unconstrained Euclidean points—fundamentally misspecify the problem by ignoring the curved structure of probability space.
Central theoretical result.
Our main contribution is establishing that calibration via Additive Log-Ratio (ALR) transformations reduces exactly to Platt scaling for binary problems (Proposition 1). This positions geometric calibration as the natural multi-class generalization of a widely-used method, providing principled justification where none previously existed. The ALR framework respects the compositional structure of probability vectors [13] while enabling standard optimization techniques in the transformed space.
@@
    Concentration (Theorem 2): Reliability scores based on Fisher-Rao distance satisfy tight concentration bounds with explicit sub-Gaussian parameter $\sigma^{2}=(1-e^{-\lambda\pi})^{2}/4$. These bounds enable sample size calculations for validation set design prior to data collection.
Geometric neutral zones.
Complementing calibration, we define reliability scores based on Fisher-Rao distance to the predicted simplex vertex (Definition 3) and construct neutral zones for principled deferral of uncertain predictions. The connection to Bhattacharyya coefficients (Remark 6) suggests potential Neyman-Pearson optimality, which we state as Conjecture 1.
Importantly, the neutral zone mechanism is method-agnostic: it can be applied to any calibration method. Our empirical analysis (Section 4.8, Table 5) confirms that operational gains are similar across calibration approaches. The contribution of geometric calibration is providing theoretical foundations, not empirical superiority.
@@
The modest scale of this application is appropriate for proof of concept; validation on larger-scale and more diverse settings remains future work (Section 5).
@@ 1.4 Novelty
Mathematical framework:
@@
Theoretical guarantees:
Theorem 2 provides the first concentration bounds for geometric reliability scores with explicit sub-Gaussian parameters. Previous work on uncertainty quantification (conformal prediction [9], deep ensembles [11]) either lacks sample complexity bounds or provides only asymptotic guarantees. Our explicit constants (Remark 10) enable practitioners to compute required validation set sizes.
Unified two-stage framework:
We decompose uncertainty-aware classification into: (i) calibration (population-level probability correction, Section 2.2) and (ii) reliability scoring (instance-level uncertainty quantification, Section 4.5). This separation clarifies that calibration alone provides modest accuracy gains (Table 4), while the combination with neutral zones (Section 4.6) achieves substantial error reduction. Our empirical analysis (Section 4.8, Table 5) demonstrates this is true regardless of calibration method, isolating the neutral zone mechanism’s contribution.
@@ 1.5 Paper Organization
Section 2 develops the mathematical framework, including geometry of the probability simplex, geometric calibration, and geometric neutral zones. Section 3 presents proven theorems and conjectures with supporting evidence. Section 4 provides empirical validation on AAV data. Section 5 discusses open problems and future directions.
@@ 3 Theoretical Results
We present three categories of results: proven theorems (Section 3.1), conjectures with strong evidence (Section 3.2), and research directions (Section 5).
@@ 3.2.1 Empirical Convergence Behavior
Validation strategy for future work: To rigorously test whether super-efficiency occurs, we propose: (i) simulation studies with known $T^{*}$ where the true calibration map is specified, (ii) computing effective dimension via $d_{\text{eff}}=\operatorname{tr}(\nabla^{2}\mathcal{L}^{*})/\|\nabla^{2}\mathcal{L}^{*}\|_{2}$, and (iii) varying regularization strength $(\lambda_{1},\lambda_{2})$ to assess persistence of the phenomenon.
@@ 3.2.2 Neyman-Pearson Optimality
Path to proof:
  1. 1.
    Binary case ($c=2$): For two classes, the Fisher-Rao distance reduces to $d_{\text{FR}}(\mathbf{p},\mathbf{e}_{1})=2\arccos(\sqrt{p_{1}})$, which is a monotone function of $p_{1}$. Under a Dirichlet noise model $\mathbf{p}^{\text{CNN}}\sim\text{Dir}(\alpha\cdot\mathbf{p}^{\text{true}})$ with concentration $\alpha>0$, the conditional density $f(d_{\text{FR}}\mid Y=1)$ versus $f(d_{\text{FR}}\mid Y=0)$ can be computed explicitly via change of variables from the Beta distribution. The likelihood ratio $f(d_{\text{FR}}\mid Y=1)/f(d_{\text{FR}}\mid Y=0)$ is then a function of $d_{\text{FR}}$ alone. We are currently deriving this density to verify monotone likelihood ratio analytically for $c=2$.
  2. 2.
[…]
  3. 3.
    General proof: The main obstacle is characterizing the distribution of $d_{\text{FR}}$ under Dirichlet perturbations on $\Delta^{c-1}$. Two potential approaches: (a) derive explicit formulas for geodesic distances under wrapped normal or Riemannian normal distributions on the simplex–these are known for certain symmetric cases [14], or (b) use concentration of measure: the positive curvature $K=+1/4$ implies Lévy-type concentration on spherical caps, which may yield sufficient control on conditional densities to establish MLR via comparison arguments. Spherical concentration is actually stronger than hyperbolic concentration for our purposes, as geodesic balls on positively curved spaces have smaller volume than Euclidean balls of the same radius.
Validation strategy: Implement Monte Carlo validation for $c\in\{2,3,4,5\}$ with varying concentration parameters $\alpha\in\{1,5,10,50\}$. For each $(c,\alpha)$ pair, generate $10^{5}$ samples and estimate likelihood ratio $\widehat{LR}(d_{\text{FR}})=\hat{f}(d_{\text{FR}}\mid\text{correct})/\hat{f}(d_{\text{FR}}\mid\text{error})$ via kernel density estimation. Check monotonicity by testing whether $\partial\widehat{LR}/\partial d_{\text{FR}}<0$ holds across the support.
@@ 4.8.3 Method Comparison
5 Open Problems and Future Directions
This section outlines research directions extending the geometric calibration framework. We organize by theme, progressing from theoretical extensions to practical questions and connections with related methods.
5.1 Curvature-Based Error Bounds
The positive curvature $K=+1/4$ of the probability simplex suggests tighter error propagation than Euclidean space. Specifically, geodesic triangles on positively curved spaces have angle sums exceeding $\pi$, and comparison theorems (Toponogov, Rauch) bound distance distortions. We conjecture:
$d_{\text{FR}}(\mathbf{p},\mathbf{q})\leq 2\tanh^{-1}\left(\frac{\|\operatorname{ALR}(\mathbf{p})-\operatorname{ALR}(\mathbf{q})\|_{2}}{C}\right)$ (18)
for some constant $C>0$, with the sublinear $\tanh^{-1}$ growth suggesting tighter bounds than linear Euclidean propagation. The intuition is that the ALR transformation “stretches” distances near the simplex boundary, while the spherical geometry “compresses” them, potentially yielding favorable cancellation.
5.2 Distribution Shift Robustness
Under covariate shift $Q\ll P$ with bounded likelihood ratio $dQ/dP\leq M$, we aim to prove:
$\left|\operatorname{\mathbb{E}}_{Q}[d_{\text{FR}}]-\operatorname{\mathbb{E}}_{P}[d_{\text{FR}}]\right|\leq\sqrt{(M-1)^{2}+2M\cdot D_{KL}(Q\|P)}\cdot\pi$ (19)
using the bounded diameter $\operatorname{diam}(\Delta^{c-1})=\pi$. This would provide guarantees for calibration transfer across related domains (e.g., training on one imaging modality, deploying on another).
5.3 Geometry vs. Temperature Scaling: Phase Transitions
Temperature scaling uses one parameter; geometric calibration uses $(c-1)^{2}+(c-1)$. We hypothesize a phase transition: geometric methods dominate when miscalibration is anisotropic, quantified by effective rank:
$r_{\text{eff}}=\frac{\operatorname{tr}(\Sigma)}{\|\Sigma\|_{2}},\quad\Sigma=\operatorname{Cov}[\operatorname{ALR}(\mathbf{p}^{\text{CNN}})].$ (20)
When $r_{\text{eff}}\approx 1$, miscalibration is nearly isotropic (all directions equally affected), and temperature scaling suffices. When $r_{\text{eff}}\gg 1$, different ALR directions require different corrections, favoring the full geometric approach. Characterizing the critical threshold $r^{*}_{\text{eff}}(c)$ as a function of number of classes remains open.
5.4 Neutral Zone Size Reduction
Does calibration reduce neutral zone size exponentially? Via Fano’s inequality, we conjecture:
$|NZ^{\text{cal}}|/|NZ^{\text{uncal}}|\stackrel{{\scriptstyle?}}{{\approx}}\exp(-\theta\cdot\Delta\text{ECE})$ (21)
for some $\theta>0$ depending on the class structure. The intuition is that improved calibration concentrates predictions away from decision boundaries, reducing the proportion requiring deferral. Empirical validation across diverse datasets would clarify the functional form.
5.5 Extensions to r-Power Methodology
The following directions extend the r-power framework [12] and will be developed in subsequent work.
5.5.1 Multi-Class Extension
Extension beyond binary classification requires testing ordered hypotheses via Union-Intersection principles. Specifically, we test:
$H_{0}^{(j)}:p_{j}\leq\max_{k\neq j}p_{k}\quad\text{vs.}\quad H_{1}^{(j)}:p_{j}>\max_{k\neq j}p_{k}$ (22)
for the predicted class $j=\hat{j}$. Our current approach uses step-down methods, adjusting for conservativeness through controlled false negative rates. The key challenge is maintaining power while controlling familywise error across the c(c1)/2c(c-1)/2 pairwise comparisons implicit in the multi-class setting.
5.5.2 ADEBO Adaptation for Bounded Support
The ADEBO estimator [24] requires approximate normality. For probability vectors with bounded support $\epsilon,1-\epsilon]^{c}\cap\Delta^{c-1}$, the logit transformation induces skewness near boundaries. Deriving moment conditions for bounded support—particularly higher-order corrections to the normal approximation—would extend ADEBO’s applicability to calibrated probability outputs.
@@ 6 Conclusion
We have developed a geometric framework for uncertainty-aware classification that bridges information geometry and statistical learning theory. The framework comprises two stages: geometric calibration of neural network probability outputs via ALR transformations, followed by reliability scoring and neutral zone construction for principled deferral of uncertain predictions.
@@
Our central theoretical result is Proposition 1, which establishes geometric calibration as the natural multi-class generalization of Platt scaling—a widely-used binary method that previously lacked principled extension to $c>2$ classes. This fills a gap in the calibration literature by grounding multi-class calibration in the intrinsic geometry of the probability simplex.
Supporting this framework, Theorem 1 provides consistency guarantees at rate $O_{p}(n^{-1/2})$ via M-estimation theory, while Theorem 2 yields concentration bounds for reliability scores with explicit sub-Gaussian parameters. These bounds enable practitioners to determine required validation set sizes before data collection—a practical benefit for applications where labeled validation data is expensive.
@@
Validation on AAV classification demonstrates that the two-stage framework achieves substantial error reduction: 72.5% error capture at 34.5% deferral rate, reducing automated decision errors from 16.8% to 6.9%. Importantly, our comparison with alternative calibration methods (Table 5) reveals that these operational gains arise primarily from the neutral zone mechanism itself, which is applicable to any well-calibrated probability output. Geometric calibration achieves competitive empirical performance while providing formal guarantees that simpler methods lack.
@@
The open problems in Section 5 outline extensions of this work. Most pressing are: (i) rigorous investigation of Conjecture 1, potentially via explicit characterization of the likelihood ratio under Dirichlet perturbations; (ii) distribution shift robustness bounds for calibration transfer across related domains; and (iii) integration with the r-power methodology [12] for distribution-free neutral zones in multi-class settings. The connection between Fisher-Rao geometry and optimal decision theory suggested by Remark 6 merits deeper exploration.
@@
As AI systems increasingly inform high-stakes decisions in medicine, criminal justice, and autonomous systems, principled uncertainty quantification becomes essential. This work contributes formal foundations for one approach to this challenge. The theoretical guarantees established here—consistency rates, concentration bounds, and the principled connection to Platt scaling—may prove particularly valuable in applications requiring rigorous validation documentation, where empirical performance benchmarks alone are insufficient.

### v2

@@ Abstract
Modern artificial intelligence systems make critical decisions yet often fail silently when uncertain—even well-calibrated models provide no mechanism to identify which specific predictions are unreliable. We develop a geometric framework addressing both calibration and instance-level uncertainty quantification for neural network probability outputs. Treating probability vectors as points on the (c1)(c-1)-dimensional probability simplex equipped with the Fisher–Rao metric, we construct: (i) Additive Log-Ratio (ALR) calibration maps that reduce exactly to Platt scaling for binary problems while extending naturally to multi-class settings, and (ii) geometric reliability scores that translate calibrated probabilities into actionable uncertainty measures, enabling principled deferral of ambiguous predictions to human review.
Theoretical contributions include: consistency of the calibration estimator at rate $O_{p}(n^{-1/2})$ via M-estimation theory (Theorem 1), and tight concentration bounds for reliability scores with explicit sub-Gaussian parameters enabling sample size calculations for validation set design (Theorem 2). We conjecture Neyman–Pearson optimality of our neutral zone construction based on connections to Bhattacharyya coefficients. Empirical validation on Adeno-Associated Virus classification demonstrates that the two-stage framework captures 72.5% of errors while deferring 34.5% of samples, reducing automated decision error rates from 16.8% to 6.9%. Notably, calibration alone yields marginal accuracy gains; the operational benefit arises primarily from the reliability scoring mechanism, which applies to any well-calibrated probability output. This work bridges information geometry and statistical learning, offering formal guarantees for uncertainty-aware classification in applications requiring rigorous validation.
@@
Preprint Note: This manuscript is under preparation for journal submission. Comments welcome at soumojit.das@wsu.edu.
@@ 1 Introduction
The fundamental issue is twofold: neural network probability outputs are miscalibrated (they do not reflect true likelihood of correctness) and they lack instance-level uncertainty quantification (they provide no mechanism to identify which specific predictions are unreliable). Guo et al. [1] demonstrated that modern deep neural networks exhibit Expected Calibration Error (ECE) often exceeding 10% on standard benchmarks. Yet even perfectly calibrated probabilities offer only population-level guarantees: knowing that 70% of predictions with confidence 0.7 are correct does not identify which 30% will fail. This gap between calibration and actionable uncertainty quantification leads to silent failures—systems appear confident while making unreliable predictions with potentially severe consequences [2, 3].
@@ 1.3 Our Approach and Contributions
a Riemannian manifold equipped with the Fisher-Rao metric (Definition 1). This geometric perspective enables a unified treatment of both calibration and uncertainty quantification, addressing the two problems identified above through a coherent mathematical framework.
Two-stage framework.
We decompose uncertainty-aware classification into complementary stages:
  1. 1.
    Calibration (population-level): Affine transformation in ALR space corrects systematic probability distortions, reducing exactly to Platt scaling for binary problems (Proposition 1) while extending naturally to multi-class settings.
  2. 2.
    Reliability scoring (instance-level): Geometric reliability scores based on Fisher-Rao distance to simplex vertices translate calibrated probabilities into actionable uncertainty measures. Neutral zones defer predictions with low reliability to human review.
This separation clarifies their distinct roles: calibration ensures probability estimates are accurate on average, while reliability scoring identifies which specific predictions warrant trust. The operational gains arise primarily from the second stage.
@@
    Concentration (Theorem 2): Reliability scores satisfy tight concentration bounds with explicit sub-Gaussian parameter $\sigma^{2}=(1-e^{-\lambda\pi})^{2}/4$. These bounds enable sample size calculations for validation set design prior to data collection—critical for applications where labeled validation data is expensive.
Method-agnostic reliability scoring.
The neutral zone mechanism can be applied to any calibration method. Our empirical analysis (Section 4.8, Table 5) confirms that operational gains are similar across calibration approaches. The contribution of geometric calibration is therefore twofold: (i) providing a principled multi-class generalization of Platt scaling grounded in information geometry, and (ii) enabling theoretical guarantees (Theorems 12) that simpler methods lack. The connection to Bhattacharyya coefficients (Remark 6) suggests potential Neyman-Pearson optimality (Conjecture 1).
@@
Notably, calibration alone improves accuracy only marginally (83.2% $\to$ 83.5%); the substantial error reduction comes from the reliability-based deferral mechanism. The modest scale of this application is appropriate for proof of concept; validation on larger-scale and more diverse settings remains future work.
@@ 1.4 Novelty
Instance-level uncertainty quantification:
While calibration methods improve probability estimates on average, they do not identify which specific predictions are reliable. Our geometric reliability scores fill this gap, translating calibrated probabilities into actionable uncertainty measures. The framework enables principled deferral: predictions with low reliability are routed to human review, while high-reliability predictions proceed automatically. Empirically, this mechanism—not calibration itself—drives the operational gains (Table 4).
Unified two-stage framework:
We decompose uncertainty-aware classification into: (i) calibration (population-level probability correction, Section 2.2) and (ii) reliability scoring (instance-level uncertainty quantification, Section 4.5). This separation clarifies that calibration alone provides modest accuracy gains, while the combination with neutral zones (Section 4.6) achieves substantial error reduction. Our empirical analysis (Section 4.8, Table 5) demonstrates this is true regardless of calibration method, isolating the neutral zone mechanism’s contribution.
Geometric foundations for multi-class calibration:
@@
Theoretical guarantees with explicit constants:
Theorem 2 provides the first concentration bounds for geometric reliability scores with explicit sub-Gaussian parameters. Previous work on uncertainty quantification (conformal prediction [9], deep ensembles [11]) either lacks sample complexity bounds or provides only asymptotic guarantees. Our explicit constants (Remark 10) enable practitioners to compute required validation set sizes before data collection.
@@ 1.5 Paper Organization
Section 2 develops the mathematical framework, including geometry of the probability simplex, geometric calibration, and geometric neutral zones. Section 3 presents proven theorems and conjectures with supporting evidence. Section 4 provides empirical validation on AAV data. Section 5 discusses limitations, future directions, and broader context.
@@ 3 Theoretical Results
We present two categories of results: proven theorems (Section 3.1) and conjectures with strong evidence (Section 3.2).
@@ 5 Conclusion
We have developed a geometric framework for uncertainty-aware classification that addresses both calibration and instance-level uncertainty quantification. The key insight is that calibration alone—even perfect calibration—provides only population-level guarantees, leaving unanswered the operationally critical question: which specific predictions should we trust? Our two-stage framework answers this by combining geometric calibration with reliability scoring that translates calibrated probabilities into actionable uncertainty measures.
@@
The primary operational contribution is the reliability scoring mechanism, which enables principled deferral of uncertain predictions. On AAV classification, this mechanism captures 72.5% of errors while deferring 34.5% of samples—performance that is largely independent of the specific calibration method used (Table 5). The geometric framework provides the theoretical foundations: Proposition 1 establishes geometric calibration as the natural multi-class generalization of Platt scaling, Theorem 1 provides consistency guarantees, and Theorem 2 yields concentration bounds with explicit sub-Gaussian parameters enabling sample size calculations before data collection.
@@
Validation on AAV classification demonstrates that the two-stage framework achieves substantial error reduction: automated decision errors drop from 16.8% to 6.9%. Crucially, calibration alone improves accuracy only marginally (83.2% $\to$ 83.5%); the operational benefit arises from reliability-based deferral. This separation clarifies that while geometric calibration provides principled theoretical foundations, the neutral zone mechanism itself—applicable to any well-calibrated output—drives the practical gains.
@@
Several open problems merit investigation: (i) rigorous proof of Conjecture 1 on Neyman-Pearson optimality; (ii) extension to high-dimensional settings ($c\gg 3$) and analysis under distribution shift; and (iii) integration with the r-power methodology [12] for distribution-free neutral zones in multi-class settings. The connection between Fisher-Rao geometry and optimal decision theory suggested by Remark 6 merits deeper exploration.
@@
As AI systems increasingly inform high-stakes decisions in medicine, criminal justice, and autonomous systems, the gap between calibration and actionable uncertainty quantification becomes critical. Well-calibrated probabilities tell us that 70% of predictions at confidence 0.7 are correct; they do not tell us which 30% will fail. This work contributes formal foundations for bridging that gap, offering instance-level reliability scores with theoretical guarantees that may prove particularly valuable in applications requiring rigorous validation documentation.
@@ A.1 Proof of Theorem 1
We start with verifying the conditions used in example 19.8 of van der Vaart.